# Sheffer polynomials and the s-ordering of exponential boson operators

### v1

@@ 3.3 Hsu–Shiue arrays and polynomials
$\mathop{{}\it HS}\nolimits_{n}(t;\,-1,1,1)=n!\,L_{n}(-t),$ $\mathop{{}\it HS}\nolimits_{n}(t;\,1,-1,-1)=(-1)^{n}n!\,L_{n}(t),$ (45)
where $L_{n}$ is the $n$’th Laguerre polynomial. By examination, negating the parameters $A,B,r$ of any exponential Riordan array of Hsu–Shiue type multiplies $\mathop{{}\it HS}\nolimits_{n,k}$ by $(-1)^{n-k}$ and alters $\mathop{{}\it HS}\nolimits_{n}(t)$ to $(-1)^{n}\mathop{{}\it HS}\nolimits_{n}(-t)$.
One can see from (39) that the polynomial sequence $n!\,L_{n}(-t)$ is the Sheffer sequence $\mathbf{S}[1/(1+\nobreak D),\allowbreak D/\allowbreak(1+\nobreak D)]$, or equivalently $\mathbf{R}[1/(1-\nobreak z),z/\allowbreak(1-\nobreak z)]$; and $(-1)^{n}n!\,L_{n}(t)$ is $\mathbf{S}[1/(1-\nobreak D),\allowbreak D/\allowbreak(1-\nobreak D)]$ or $\mathbf{R}[1/(1+\nobreak z),z/\allowbreak(1+\nobreak z)]$. (Cf. [9].)
@@ 4 Exponentiated operators
The main result of this paper, on the $s$-ordered representation of the exponential of a single-mode “boson string” of the form ${a^{\dagger}}^{L}a{a^{\dagger}}^{R}$, will be deduced in Sec. 5 from a restricted result on a normal ordering; or equally well, an anti-normal one. The normal and anti-normal results, given in Theorem 4.4 below, appeared in the Introduction as eqs. (3),(4). They will follow from Theorem 4.2, a result of the author on the $n$’th power of ${a^{\dagger}}^{L}a{a^{\dagger}}^{R}$ [41, Thm. 6.5], a short proof of which is supplied, to make this paper self-contained.
As will be sketched, they also follow from the results of Lang [14, 15] and Błasiak et al. [16, 17] on exponentiated vector fields and on applications of Sheffer sequences to ordering problems. That Fan et al. [42] obtained the $R=\nobreak 0$ case of eq. (3) by the formal IWOP (integration within ordered products) technique [43] must also be mentioned.
As motivation for introducing the coefficients of Hsu–Shiue and other Sheffer polynomials, consider the following identities. Firstly, the two in $\mathcal{W}[[\lambda]]$,
${\rm e}^{\lambda({a^{\dagger}}a)}$ $=\mathopen{:}\,\mathrel{\exp\left[\left({\rm e}^{\lambda}-1\right){a^{\dagger}}a\right]}\,\mathclose{:}_{N}$ (46a)
$=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda;\,0,1,0)}\,\mathclose{:}_{N}=\mathopen{:}\,\mathrel{\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,0,1,0)\frac{\lambda^{n}}{n!}}\,\mathclose{:}_{N}$ (46b)
@@
${\rm e}^{\lambda(a{a^{\dagger}})}$ $=\mathopen{:}\,\mathrel{{\rm e}^{\lambda}\exp\left[\left({\rm e}^{\lambda}-1\right)a{a^{\dagger}}\right]}\,\mathclose{:}_{N}$ (46a)
$=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda;\,0,1,1)}\,\mathclose{:}_{N}=\mathopen{:}\,\mathrel{\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,0,1,1)\frac{\lambda^{n}}{n!}}\,\mathclose{:}_{N}.$ (46b)
Here, (46a) is the Cahill–Glauber normal ordering shown in (1), and (46a) is an immediate consequence. The rewritings in terms of $\mathcal{HS}(t,z)$ come from the $A\to 0$ limit of the EGF formula in Proposition 3.4. Secondly, the two in $\mathcal{W}$,
$({a^{\dagger}}a)^{n}$ $=\sum_{k=0}^{n}\left\{{{n}\atop{k}}\right\}{a^{\dagger}}^{k}a^{k}=\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(0,1,0)\,{a^{\dagger}}^{k}a^{k}$ (47)
@@
$(a{a^{\dagger}})^{n}$ $=\sum_{k=0}^{n}\left\{{{n+1}\atop{k+1}}\right\}{a^{\dagger}}^{k}a^{k}=\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(0,1,1)\,{a^{\dagger}}^{k}a^{k}.$ (48)
These two are due to Katriel [44], who showed for the first time how Stirling numbers arise in boson algebra. Each can also be proved combinatorially, by applying Wick’s theorem and enumerating contractions [45, §9.3].
Clearly (46b),(46b) follow from (47),(48) by multiplying by $\lambda^{n}/n!$ and summing over $n$. What is needed for a generalization to the exponential operator ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ is a formula for $({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}$. Duchamp et al. proved the following.
@@
$\omega^{n}=\left({a^{\dagger}}^{e}\right)^{n}\left[\sum_{k=0}^{n}\mathbf{R}[d,h]_{n,k}\,{a^{\dagger}}^{k}a^{k}\right],$ (49)
@@
What is difficult is finding explicitly the $(L,R)$-dependent series $d,h\in\mathbb{C}[[z]]$ that define the Riordan array. Identities (46a),(46b) suggest that $\mathbf{R}[d,h]$ should be of Hsu–Shiue type. But the derivation of the identity (49) need not be combinatorial, or based on a generalization of Wick’s theorem.
Combinatorial interpretations of the reordered coefficients of such boson strings as ${a^{\dagger}}^{r_{n}}a^{s_{n}}\dots a^{\dagger}{}^{r_{1}}a^{s_{1}}\in\mathcal{W}$ are known [18]. Also, interpretations of the coefficients $\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)$ for general integer values of $A,B,r$ are known, such as ones involving the placing of balls in cells and their compartments, or the drawing of balls from an urn. But such combinatorial interpretations tend at present to be complicated. (See, e.g., [35, Chapter 4].) The following two parametric identities in $\mathcal{W}$ are extracted from [41, Theorem 6.5], and have a non-combinatorial proof. The first subsumes Katriel’s formulae, (47) and (48).
@@
$\begin{aligned} \left({a^{\dagger}}^{L}a{a^{\dagger}}^{R}\right)^{n}&=\left({a^{\dagger}}^{e}\right)^{n}\left[\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(-e,1,R)\,{a^{\dagger}}^{k}a^{k}\right]\\ &=\mathopen{:}\,\mathrel{\left({a^{\dagger}}^{e}\right)^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,-e,1,R)}\,\mathclose{:}_{N}\end{aligned}$ (50)
@@
$\begin{aligned} \left({a^{\dagger}}^{L}a{a^{\dagger}}^{R}\right)^{n}&=\left[\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(e,-1,-L)\,a^{k}{a^{\dagger}}^{k}\right]\left({a^{\dagger}}^{e}\right)^{n}\\ &=\mathopen{:}\,\mathrel{\mathop{{}\it HS}\nolimits_{n}(a{a^{\dagger}};\,e,-1,-L)\left({a^{\dagger}}^{e}\right)^{n}}\,\mathclose{:}_{A},\end{aligned}$ (51)
where the $\mathop{{}\it HS}\nolimits_{n,k}$ coefficients can be computed from the summation formula (40) or the recurrence (42).
@@
Representing $a,{a^{\dagger}}$ by $D_{x},x$, acting on any convenient space of functions of $x$, yields a faithful representation of $\mathcal{W}$, so it suffices to obtain versions of (50),(51) with $a,{a^{\dagger}}$ replaced by $D_{x},x$.
Replacing $x$ by $xD_{x}$ in the definition (34) of the coefficients $\mathop{{}\it HS}\nolimits_{n,k}$ yields
$(xD_{x}+r)^{\underline{n},A}=\sum_{k=0}^{n}HS_{n,k}(A,B,r)\,(xD_{x})^{\underline{k},B}.$ (52)
@@
$(xD_{x}+r)^{\underline{n},A}$ $=x^{-r}\left[(x^{A})^{n}(x^{1-A}D_{x})^{n}\right]x^{r}=(x^{A})^{n}(x^{1-A-r}D_{x}x^{r})^{n},$ (54a)
$=x^{-r}\left[(xD_{x}x^{A})^{n}(x^{-A})^{n}\right]x^{r}=(x^{1-r}D_{x}x^{A+r})^{n}(x^{-A})^{n}$ (54b)
@@
To obtain a version of (50) with $a,{a^{\dagger}}$ replaced by $D_{x},x$, rewrite the left-hand side of (52) with the aid of (54a), and the right-hand side with the aid of (53a); and also set (A,B,r)=(e,1,R)(A,B,r)=(-e,1,R). To obtain (51) similarly, use (54b) and (53b), and set $(A,B,r)=\allowbreak(e,-1,1-\nobreak L)$. ∎
@@
If $L=R=1$, identities (50),(51) of the theorem reduce to
$\left({a^{\dagger}}a{a^{\dagger}}\right)^{n}=\left\{\begin{aligned} &\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,-1,1,1)}\,\mathclose{:}_{N}=n!\;\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}L_{n}(-{a^{\dagger}}a)}\,\mathclose{:}_{N},\\ &\mathopen{:}\,\mathrel{\mathop{{}\it HS}\nolimits_{n}(a{a^{\dagger}};\,1,-1,-1)a^{n}}\,\mathclose{:}_{A}=(-1)^{n}n!\;\mathopen{:}\,\mathrel{L_{n}(a{a^{\dagger}}){a^{\dagger}}^{n}}\,\mathclose{:}_{A},\end{aligned}\right$. (55)
by (45), so that
$\left({a^{\dagger}}a{a^{\dagger}}\right)^{n}=\left\{\begin{aligned} &{a^{\dagger}}^{n}\cdot\sum_{k=0}^{n}\frac{n!^{2}}{k!^{2}(n-k)!}\,{a^{\dagger}}^{k}a^{k},\\ &\sum_{k=0}^{n}(-1)^{n-k}\frac{n!^{2}}{k!^{2}(n-k)!}\,a^{k}{a^{\dagger}}^{k}\cdot{a^{\dagger}}^{n}.\end{aligned}\right$. (56)
These are known Laguerre-polynomial representations [48].
Multiplying the identities in Theorem 4.2 by $\lambda^{n}/n!$ and summing over $n$ yields the following identities in $\mathcal{W}[[\lambda]]$, with the two explicit expressions coming from the EGF formula in Proposition 3.4. The first subsumes (46b) and (46b).
@@
$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda{a^{\dagger}}^{e};\,-e,1,R)}\,\mathclose{:}_{N}\\ &=\mathopen{:}\,\mathrel{(1-e\lambda{a^{\dagger}}^{e})^{-R/e}\exp\left\{\left[\bigl{(}1-e\lambda{a^{\dagger}}^{e}\bigr{)}^{-1/e}-1\right]{a^{\dagger}}a\right\}}\,\mathclose{:}_{N}\end{aligned}$ (57)
@@
$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}(a{a^{\dagger}},\lambda{a^{\dagger}}^{e};\,e,-1,-L)}\,\mathclose{:}_{A}\\ &=\mathopen{:}\,\mathrel{\exp\left\{-a{a^{\dagger}}\left[\bigl{(}1+e\lambda{a^{\dagger}}^{e}\bigr{)}^{-1/e}-1\right]\right\}(1+e\lambda{a^{\dagger}}^{e})^{-L/e}}\,\mathclose{:}_{A},\end{aligned}$ (58)
@@
The $R=0$ case of (57) was previously derived by Lang [14, 15] and Błasiak et al. [16]. They used a representation of $\mathcal{W}$ in which $a,{a^{\dagger}}$ are represented by $D_{x},x$, and exponentiated ${a^{\dagger}}^{L}a$ by integrating the vector field $x^{L}D_{x}$. Also, Penson et al. [48] obtained what amounted to the $R=1$ case of (57) by exploiting a Dobiński representation of generalized Bell polynomials.
More generally, Błasiak et al. [16] integrated $q(x)D_{x}+v(x)$, where $q,v\in\mathbb{C}[[x]]$. As ${a^{\dagger}}^{L}a{a^{\dagger}}^{R}$ equals $({a^{\dagger}})^{L+R}a+\nobreak R({a^{\dagger}})^{L+R-1}$, by substituting $q(x)=x^{L+R}$ and $v(x)=Rx^{L+R-1}$ into their formulae one can derive (57) in full generality.
@@
$\begin{split}&\exp\left\{\lambda\,\frac{1}{f^{\prime}({a^{\dagger}})}\left[a-\frac{g^{\prime}({a^{\dagger}})}{g({a^{\dagger}})}\right]\right\}\\ &\quad=\mathopen{:}\,\mathrel{\frac{g({a^{\dagger}})}{g\left(\bar{f}(f({a^{\dagger}})+\lambda)\right)}\exp{\Bigl{\{}\left[\bar{f}(f({a^{\dagger}})+\lambda)-{a^{\dagger}}\right]a\Bigr{\}}}}\,\mathclose{:}_{N},\end{split}$ (59)
where $g,f\in\mathbb{C}[[D]]$ are any pair that defines a Sheffer sequence. Equation (59) was not obtained in the power-series framework of the present paper, as its derivation employed a Fock space representation and projections onto coherent states, but it is a valid identity in $\mathbb{C}[[{a^{\dagger}}]][a][[\lambda]]$.
By examination, formally substituting $g(D)=D^{-R}$, $f(D)=-D^{-e}/e$ into (59) yields (57). But this additional derivation of eq. (57) is not rigorous, as these $g,f$ contain negative powers of $D$ and do not belong to $\mathbb{C}[[D]]$. At most, it is an analytic continuation of a rigorous derivation.

### v2

@@ 3.3 Hsu–Shiue arrays and polynomials
$\mathop{{}\it HS}\nolimits_{n}(t;\,-1,1,1)=n!\,L_{n}(-t),$ $\mathop{{}\it HS}\nolimits_{n}(t;\,1,-1,-1)=(-1)^{n}n!\,L_{n}(t),$ (3.22)
where $L_{n}$ is the $n$’th Laguerre polynomial.
By examination, multiplying the parameters $A,B,r$ of any exponential Riordan array $\mathcal{HS}=(\mathop{{}\it HS}\nolimits_{n,k})$ of Hsu–Shiue type by $c\neq 0$ multiplies $\mathop{{}\it HS}\nolimits_{n,k}$ by $c^{n-k}$ and alters $\mathop{{}\it HS}\nolimits_{n}(t)$ to $c^{n}\mathop{{}\it HS}\nolimits_{n}(t/c)$. As an application, the $m=n$ case of the $s$-ordering conversion formula (2.13) can be rewritten as
$\mathopen{:}\,\mathrel{({a^{\dagger}}a)^{n}}\,\mathclose{:}_{s}=\mathopen{:}\,\mathrel{\delta^{n}n!\,L_{n}(-{a^{\dagger}}a/\delta)}\,\mathclose{:}_{s^{\prime}}=\mathopen{:}\,\mathrel{\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,-\delta,\delta,\delta)}\,\mathclose{:}_{s^{\prime}},$ (3.23)
where $\delta=(s-s^{\prime})/2$. The arrays $\mathcal{HS}(-\delta,\delta,\delta)$, $\delta\in\mathbb{C}$, make up an ‘$s$-shifting’ subgroup of $\mathfrak{R}$ that is isomorphic to the additive group of $\mathbb{C}$.
One can see from (3.16) that the polynomial sequence $n!\,L_{n}(-t)$ is the Sheffer sequence $\mathbf{S}[1/(1+\penalty 10000D),\allowbreak D/\allowbreak(1+\penalty 10000D)]$, or equivalently $\mathbf{R}[1/(1-\penalty 10000z),z/\allowbreak(1-\penalty 10000z)]$; and $(-1)^{n}n!\,L_{n}(t)$ is $\mathbf{S}[1/(1-\penalty 10000D),\allowbreak D/\allowbreak(1-\penalty 10000D)]$ or $\mathbf{R}[1/(1+\penalty 10000z),z/\allowbreak(1+\penalty 10000z)]$. (Cf. [9].) The just-mentioned $1$-parameter subgroup of $\mathfrak{R}$ comprises $\mathbf{S}[1/(1+\penalty 10000\delta D],\allowbreak D/(1+\penalty 10000\delta D)]$ or $\allowbreak\mathbf{R}[1/(1-\penalty 10000\delta z),\allowbreak z/(1-\penalty 10000\delta z)]$, $\delta\in\mathbb{C}$.
@@ 4 Exponentiated operators
The main result of this paper, on the $s$-ordered representation of the exponential of a single-mode “boson string” of the form ${a^{\dagger}}^{L}a{a^{\dagger}}^{R}$, will be deduced in Sec. 5 from a restricted result on a normal ordering; or equally well, an anti-normal one. The normal and anti-normal results, given in Theorem 4.4 below, appeared in the Introduction as eqs. (1.4),(1.5). They will follow from Theorem 4.2, a result of the author on the $n$’th power of ${a^{\dagger}}^{L}a{a^{\dagger}}^{R}$ [41, Thm. 6.5], a short proof of which is supplied, to make this paper self-contained.
[…]
To introduce further the coefficients of Hsu–Shiue and other Sheffer polynomials, consider the following identities. Firstly, the two in $\mathcal{W}[[\lambda]]$,
${\rm e}^{\lambda({a^{\dagger}}a)}$ $=\mathopen{:}\,\mathrel{\exp\left[\left({\rm e}^{\lambda}-1\right){a^{\dagger}}a\right]}\,\mathclose{:}_{N}$ (4.1a)
$=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda;\,0,1,0)}\,\mathclose{:}_{N}=\mathopen{:}\,\mathrel{\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,0,1,0)\frac{\lambda^{n}}{n!}}\,\mathclose{:}_{N}$ (4.1b)
@@
${\rm e}^{\lambda(a{a^{\dagger}})}$ $=\mathopen{:}\,\mathrel{{\rm e}^{\lambda}\exp\left[\left({\rm e}^{\lambda}-1\right)a{a^{\dagger}}\right]}\,\mathclose{:}_{N}$ (4.1a)
$=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda;\,0,1,1)}\,\mathclose{:}_{N}=\mathopen{:}\,\mathrel{\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,0,1,1)\frac{\lambda^{n}}{n!}}\,\mathclose{:}_{N}.$ (4.1b)
Here, (4.1a) is the Cahill–Glauber normal ordering shown in (1.1), and (4.1a) is an immediate consequence. The rewritings in terms of $\mathcal{HS}(t,z)$ come from the $A\to 0$ limit of the EGF formula in Proposition 3.4. Secondly, the two in $\mathcal{W}$,
$({a^{\dagger}}a)^{n}$ $=\sum_{k=0}^{n}\left\{{{n}\atop{k}}\right\}{a^{\dagger}}^{k}a^{k}=\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(0,1,0)\,{a^{\dagger}}^{k}a^{k}$ (4.2)
@@
$(a{a^{\dagger}})^{n}$ $=\sum_{k=0}^{n}\left\{{{n+1}\atop{k+1}}\right\}{a^{\dagger}}^{k}a^{k}=\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(0,1,1)\,{a^{\dagger}}^{k}a^{k}.$ (4.3)
These are due to Katriel [44], who showed for the first time how Stirling numbers arise in boson algebra. Each can also be proved combinatorially, by applying Wick’s theorem and enumerating contractions [45, §9.3]. More generally, one can expand any polynomial in ${a^{\dagger}}a$ or $a{a^{\dagger}}$ in the monomials ${a^{\dagger}}^{k}a^{k}$ or $a^{k}{a^{\dagger}}^{k}$ [46].
Clearly (4.1b),(4.1b) follow from (4.2),(4.3) by multiplying by $\lambda^{n}/n!$ and summing over $n$. What is needed for a generalization to the exponential operator ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ is a formula for $({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}$. Duchamp et al. proved the following.
@@
$\omega^{n}=\left({a^{\dagger}}^{e}\right)^{n}\left[\sum_{k=0}^{n}\mathbf{R}[d,h]_{n,k}\,{a^{\dagger}}^{k}a^{k}\right],$ (4.4)
@@
What is difficult is finding explicitly the $(L,R)$-dependent series $d,h\in\mathbb{C}[[z]]$ that define the Riordan array. Identities (4.1a),(4.1b) suggest that $\mathbf{R}[d,h]$ should be of Hsu–Shiue type. But the derivation of the identity (4.4) need not be combinatorial, or based on a generalization of Wick’s theorem.
Combinatorial interpretations of the reordered coefficients of such boson strings as ${a^{\dagger}}^{r_{n}}a^{s_{n}}\dots a^{\dagger}{}^{r_{1}}a^{s_{1}}\in\mathcal{W}$ are known [18]. Also, interpretations of the coefficients $\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)$ for general integer values of $A,B,r$ are known, such as ones involving the placing of balls in cells and their compartments, or the drawing of balls from an urn. But such combinatorial interpretations tend at present to be complicated. (See, e.g., [35, Chapter 4].) The following two parametric identities in $\mathcal{W}$ are extracted from [41, Theorem 6.5], and have a non-combinatorial proof. The first subsumes Katriel’s formulae, (4.2) and (4.3).
@@
$\begin{aligned} \left({a^{\dagger}}^{L}a{a^{\dagger}}^{R}\right)^{n}&=\left({a^{\dagger}}^{e}\right)^{n}\left[\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(-e,1,R)\,{a^{\dagger}}^{k}a^{k}\right]\\ &=\mathopen{:}\,\mathrel{\left({a^{\dagger}}^{e}\right)^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,-e,1,R)}\,\mathclose{:}_{N}\end{aligned}$ (4.5)
@@
$\begin{aligned} \left({a^{\dagger}}^{L}a{a^{\dagger}}^{R}\right)^{n}&=\left[\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(e,-1,-L)\,a^{k}{a^{\dagger}}^{k}\right]\left({a^{\dagger}}^{e}\right)^{n}\\ &=\mathopen{:}\,\mathrel{\mathop{{}\it HS}\nolimits_{n}(a{a^{\dagger}};\,e,-1,-L)\left({a^{\dagger}}^{e}\right)^{n}}\,\mathclose{:}_{A},\end{aligned}$ (4.6)
where the $\mathop{{}\it HS}\nolimits_{n,k}$ coefficients can be computed from the summation formula (3.17) or the recurrence (3.19).
@@
Representing $a,{a^{\dagger}}$ by $D_{x},x$, acting on any convenient space of functions of $x$, yields a faithful representation of $\mathcal{W}$, so it suffices to obtain versions of (4.5),(4.6) with $a,{a^{\dagger}}$ replaced by $D_{x},x$.
Replacing $x$ by $xD_{x}$ in the definition (3.11) of the coefficients $\mathop{{}\it HS}\nolimits_{n,k}$ yields
$(xD_{x}+r)^{\underline{n},A}=\sum_{k=0}^{n}HS_{n,k}(A,B,r)\,(xD_{x})^{\underline{k},B}.$ (4.7)
@@
$(xD_{x}+r)^{\underline{n},A}$ $=x^{-r}\left[(x^{A})^{n}(x^{1-A}D_{x})^{n}\right]x^{r}=(x^{A})^{n}(x^{1-A-r}D_{x}x^{r})^{n},$ (4.9a)
$=x^{-r}\left[(xD_{x}x^{A})^{n}(x^{-A})^{n}\right]x^{r}=(x^{1-r}D_{x}x^{A+r})^{n}(x^{-A})^{n}$ (4.9b)
@@
To obtain a version of (4.5) with $a,{a^{\dagger}}$ replaced by $D_{x},x$, rewrite the left-hand side of (4.7) with the aid of (4.9a), and the right-hand side with the aid of (4.8a); and also set (A,B,r)=(e,1,R)(A,B,r)=(-e,1,R). To obtain (4.6) similarly, use (4.9b) and (4.8b), and set $(A,B,r)=\allowbreak(e,-1,1-\penalty 10000L)$. ∎
@@
If $L=R=1$, identities (4.5),(4.6) of the theorem reduce to
$\left({a^{\dagger}}a{a^{\dagger}}\right)^{n}=\left\{\begin{aligned} &\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,-1,1,1)}\,\mathclose{:}_{N}=n!\;\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}L_{n}(-{a^{\dagger}}a)}\,\mathclose{:}_{N},\\ &\mathopen{:}\,\mathrel{\mathop{{}\it HS}\nolimits_{n}(a{a^{\dagger}};\,1,-1,-1)a^{n}}\,\mathclose{:}_{A}=(-1)^{n}n!\;\mathopen{:}\,\mathrel{L_{n}(a{a^{\dagger}}){a^{\dagger}}^{n}}\,\mathclose{:}_{A},\end{aligned}\right$. (4.10)
by (3.22), so that
$\left({a^{\dagger}}a{a^{\dagger}}\right)^{n}=\left\{\begin{aligned} &{a^{\dagger}}^{n}\cdot\sum_{k=0}^{n}\frac{n!^{2}}{k!^{2}(n-k)!}\,{a^{\dagger}}^{k}a^{k},\\ &\sum_{k=0}^{n}(-1)^{n-k}\frac{n!^{2}}{k!^{2}(n-k)!}\,a^{k}{a^{\dagger}}^{k}\cdot{a^{\dagger}}^{n}.\end{aligned}\right$. (4.11)
These are known Laguerre-polynomial representations [49], which by the $(s,s^{\prime})=(\pm 1,\mp 1)$ cases of the conversion formula (2.13) can be seen to be equivalent to $({a^{\dagger}}a{a^{\dagger}})^{n}={a^{\dagger}}^{n}a^{n}{a^{\dagger}}^{n}$, the Viskov identity [35].
Multiplying the identities in Theorem 4.2 by $\lambda^{n}/n!$ and summing over $n$ yields the following identities in $\mathcal{W}[[\lambda]]$, with the two explicit expressions coming from the EGF formula in Proposition 3.4. The first subsumes (4.1b) and (4.1b).
@@
$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda{a^{\dagger}}^{e};\,-e,1,R)}\,\mathclose{:}_{N}\\ &=\mathopen{:}\,\mathrel{(1-e\lambda{a^{\dagger}}^{e})^{-R/e}\exp\left\{\left[\bigl(1-e\lambda{a^{\dagger}}^{e}\bigr)^{-1/e}-1\right]{a^{\dagger}}a\right\}}\,\mathclose{:}_{N}\end{aligned}$ (4.12)
@@
$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}(a{a^{\dagger}},\lambda{a^{\dagger}}^{e};\,e,-1,-L)}\,\mathclose{:}_{A}\\ &=\mathopen{:}\,\mathrel{\exp\left\{-a{a^{\dagger}}\left[\bigl(1+e\lambda{a^{\dagger}}^{e}\bigr)^{-1/e}-1\right]\right\}(1+e\lambda{a^{\dagger}}^{e})^{-L/e}}\,\mathclose{:}_{A},\end{aligned}$ (4.13)
@@
The $R=0$ case of (4.12) was previously derived by Lang [14, 15] and Błasiak et al. [16]. They used a representation of $\mathcal{W}$ in which $a,{a^{\dagger}}$ are represented by $D_{x},x$, and exponentiated ${a^{\dagger}}^{L}a$ by integrating the vector field $x^{L}D_{x}$. Also, Penson et al. [49] obtained what amounted to the $R=1$ case of (4.12) by exploiting a Dobiński representation of generalized Bell polynomials.
More generally, Błasiak et al. [16] integrated $q(x)D_{x}+v(x)$, where $q,v\in\mathbb{C}[[x]]$. As ${a^{\dagger}}^{L}a{a^{\dagger}}^{R}$ equals $({a^{\dagger}})^{L+R}a+\penalty 10000R({a^{\dagger}})^{L+R-1}$, by substituting $q(x)=x^{L+R}$ and $v(x)=Rx^{L+R-1}$ into their formulae one can derive (4.12) in full generality.
@@
$\begin{split}&\exp\left\{\lambda\,\frac{1}{f^{\prime}({a^{\dagger}})}\left[a-\frac{g^{\prime}({a^{\dagger}})}{g({a^{\dagger}})}\right]\right\}\\ &\quad=\mathopen{:}\,\mathrel{\frac{g({a^{\dagger}})}{g\left(\bar{f}(f({a^{\dagger}})+\lambda)\right)}\exp{\Bigl\{\left[\bar{f}(f({a^{\dagger}})+\lambda)-{a^{\dagger}}\right]a\Bigr\}}}\,\mathclose{:}_{N},\end{split}$ (4.14)
where $g,f\in\mathbb{C}[[D]]$ are any pair that defines a Sheffer sequence. Equation (4.14) was not obtained in the power-series framework of the present paper, as its derivation employed a Fock space representation and projections onto coherent states, but it is a valid identity in $\mathbb{C}[[{a^{\dagger}}]][a][[\lambda]]$.
By examination, formally substituting $g(D)=D^{-R}$, $f(D)=-D^{-e}/e$ into (4.14) yields (4.12). But this additional derivation of eq. (4.12) is not rigorous, as these $g,f$ contain negative powers of $D$ and do not belong to $\mathbb{C}[[D]]$. At most, it is an analytic continuation of a rigorous derivation.